intel) LEIL ofind allthe sked and evenbally ved part d the Dis.
Answer #1

x=1/2 and x=3/2 are fixed and eventually fixed points respectively.We have to find all fixed and eventually fixed point of os fui= 10-11 e in R. solution To find fixed point ? Ret fao-x=0 = 12We can clearl eck . that Clearly from second condition, we can say that Yo is fixed point- and the only fixed point of few=pe